[Hallc_running] [New Logentry] Hall C RC update: 10/23/21

mack at jlab.org mack at jlab.org
Sat Oct 23 09:50:03 EDT 2021

Logentry Text:
Trippy beam came back Friday evening. It was possible to check the beam spot size and beam vs target alignment, but there was no significant production overnite. The beam current was briefly reduced to Halls A and C at one point to see if that would reduce trips. 

>From the PD update this morning, a major source of problems was BLM trips near the now unpowered magnet in Arc 7. It sounds like the stray field on the adjacent arc had dramatically changed, so the thought is that the operators have to steer around that. 

As of 08:48, the shift crew reported that Hall C is back to 42 muA. 
Hall A pass change still tentatively Sunday. Secretary of Energy tour on Monday morning.

The running total of pions is essentially same as 24 hours ago. There were no good runs. 

We will have a brief RC meeting Friday at 10am. I don't see any point in rending our garments over the lost production, but there might be something constructive you'd like me to follow up on.

Bluejeans link: https://bluejeans.com/861439920
RC meeting page (stale page. I'll update it later for posterity): https://hallcweb.jlab.org/wiki/index.php/RC_Meeting,_Friday,_October_22_,_2021


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3930959
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211023/be751619/attachment.html>

More information about the Hallc_running mailing list